OpenStudy (anonymous):

A circle has a radius of 1.5 m. Which the angle subtended at the center of the circle by an arc of length 1m?

OpenStudy (jdoe0001):

\(\bf s = r\theta\implies \cfrac{s}{r}=\theta \\ \quad \\ r=radius\qquad s=\textit{arc length}\qquad \theta = \textit{central angle}\)

OpenStudy (anonymous):

s = r *theta s = arc length r = radius theta = central angle 1m = 1.5m*theta theta = 1/1.5m theta = ?

OpenStudy (anonymous):

2/3 radians - convert that into degrees that's your answer

OpenStudy (anonymous):

Still not managed to solve this problem.The circle has a radius of 1.5 m. Which the angle subtended at the center of the circle by an arc of length 1m?

OpenStudy (anonymous):

It's all explained there..all you have to do is convert 2/3 radians into degrees...